Capricorn77:

Solve the equation and check the solution. a minus 2 and one-half = 1 and one-half

Laylalyssa:

Are there any choices orrrrrrrr

supie:

...orrrrrr XD

Capricorn77:

no there isn't your supposed to answer it

jacky66:

right??

supie:

I thought there was answer choices.......

supie:

but ok

jacky66:

supie

supie:

yes?

Capricorn77:

ok this isn't supposed to be a chat

supie:

We arent chatting

jacky66:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie yes? \(\color{#0cbb34}{\text{End of Quote}}\) nun'it's just catchie

supie:

lol

jhonyy9:

\[a -2\frac{ 1}{ 2 } = 1\frac{ 1 }{ 2}\] so this is your equation what you need to solve it for ,,a"

Capricorn77:

ok